Operations Report – April 20th

Crew 245 Operations Report 20-04-2021

SOL: 10

Name of person filing report: Shravan Hariharan

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator: Working nominally

Hours run: 7.5

From what time last night: 9:00 pm, SOC 63%

To what time this morning: 4:30 am, SOC 100%

List any additional daytime hours when the generator was run: N/A

Solar— SOC 73% at 6:58 pm (Before generator is run at night)

Notes on power system: For the last several sols, the freezers in the Science Dome have been running in order to preserve science samples. Due to this, the station power consumption has been significantly higher than it previously was, resulting in the generator being turned on earlier (as early as 7:00 pm on colder days, and 9:00 pm on warmer days) once the SOC reaches below 70%.

Diesel Reading – 50%

Station Propane Reading – 70%

Water (loft tank): 40 gallons

Water Meter: 150619.5 units

Water (static tank): 175 gallons – estimated from checking remaining static tank volume, confirmation from Outpost would be appreciated!

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Spirit rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Opportunity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: no, in town

Curiosity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Perseverance rover used: No

Hours: 204.1

Beginning charge: 100%

Ending charge: 84%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: 350.1

Reason for use: EVA

Oil Added? No

ATV Fuel Used: 1 gallon (estimated)

# Hours the ATVs were used today: 4 (but 1 hour of active transit)

Notes on ATVs: Mission Director will monitor and refuel ATVs for duration of mission.

HabCar used and why, where?: Nothing to report

CrewCar used and why, where?: Nothing to report

General notes and comments: Nothing to report

Summary of internet: Internet is working nominally. Crew is rationing internet usage during the day to ensure that there is sufficient bandwidth for Mission Support Comms window.

Summary of suits and radios: All radios nominal, all suits currently charging. Suits 3, 6, and 7 used on EVA today and all functioned nominally.

Summary of Hab operations: Nothing to report besides nominal crew activities.

Summary of GreenHab operations: Being used to host Crew Botanist botany experiment. Crew Botanist briefly checked plant growth and health in GreenHab today.

Summary of ScienceDome operations: Used by Crew Scientist and HSO to monitor HSO botany experiment, and used by Crew Scientist and XO to conduct soil sample processing.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: To Outpost – can our estimate of our remaining water be checked? We opened up the static tank to check the volume, but prior to planning out our remaining water usage, a second pair of eyes would be greatly appreciated!

-Red Planet People – MDRS Crew 245 "Team Patamars"To Mars and Beyond – For All!